OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 2 > -10 and 2x less than or greater to 12 A.) A number line with a closed circle on -8, an open circle on 6, and shading in between. B.) A number line with a closed circle on -8, shading to the left, and an open circle on 6, shading to the right. C.) A number line with an open circle on -8, shading to the left, and a closed circle on 6, shading to the right. D.) A number line with an open circle on -8, a closed circle on 6, and shading in between.

OpenStudy (anonymous):

hold the phone there is no such thing as "less than or greater to" maybe "less than or equal to"?

OpenStudy (anonymous):

\(x-2>-10\) or \(2x\leq 12\) ?

OpenStudy (anonymous):

we can solve each separately \[x-2>-10\] \[x>-8\] OR \[2x\leq 12\] \[x\leq 6\]

OpenStudy (anonymous):

oops should be "AND" in which case it would be \[-8<x\leq 6\]

OpenStudy (anonymous):

so picture would look like A number line with an open circle on -8, a closed circle on 6, and shading in between.

OpenStudy (anonymous):

Thanks for you help :)

OpenStudy (anonymous):

yw